# RAG-Optimized Tibetan Tourism LLMs: Enhancing Accuracy and Personalization

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: September 20-22, 2024; Xiamen, China; isbn: 979-8-4007-1717-8submissionid: XA5000ccs: Computing methodologies Natural language generationccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processing
@@ 1.1. Current issues regarding Tibet’s cultural tourism and LLMs
To address these challenges, this study, under Project ID: XZ2024-01ZY0008, titled ”Research and Application of Intelligent Tourism Service System in Tibet Based on LLM,” aims to enhance tourists’ travel experiences and promote the development of Tibet’s tourism industry. The project focuses on the systematic organization and classification of viewpoints in the region. This includes detailed documentation of each viewpoint’s historical background, geographical location, cultural significance, and transportation information (Chen2020).
@@
Regarding the insufficiency in personalized recommendation capabilities, traditional models emphasize generalization but lack targeted fine-tuning, leading to suboptimal performance in personalized tourism contexts. Specifically, these models may exhibit ”hallucination” phenomena(Chen2023), where they fail to accurately understand user needs and recommend the most suitable attractions. This limitation is particularly evident in the context of Tibetan tourism, where the unique characteristics of many emerging attractions are not effectively captured, resulting in recommendations that lack accuracy and relevance.
Our research addresses these challenges by employing a vector database query method within the Retrieval-Augmented Generation (RAG) framework(meta_rag_2020). LLM extracts user requirements, which are then matched against a vector database to identify suitable attractions. These viewpoints are returned to the LLM for integration, resulting in accurate and contextually relevant recommendations tailored to the user’s preferences.
@@ 2.1. Vectorization Methods: TF-IDF and BERT
TF-IDF (Term Frequency-Inverse Document Frequency) is used to evaluate the importance of a word in a document collection or corpus. By multiplying TF (term frequency) and IDF (inverse document frequency), the TF-IDF value is obtained, which measures the significance of a word to a particular document(Havrlant2017).
BERT (Bidirectional Encoder Representations from Transformers) handles unlabelled text data by pre-training deep bidirectional representations(Devlin2019). This pre-training process involves jointly conditioning on both left and right context in all layers, allowing the pre-trained BERT model to be fine-tuned for various tasks by simply adding an output layer. The advantage of this approach lies in its flexibility to adapt to the text data of this project and its capability to create state-of-the-art models for a wide range of NLP tasks.
@@ 2.2. Retrieval Construction Methods
Flat (Brute Force Search): The system compares each vector in the database with the query vector, calculates their distance or similarity, and returns the closest vectors(Li2020).
HNSW (Hierarchical Navigable Small World): This method constructs a layered, small-world graph structure to accelerate query speed. HNSW supports multiple distance metrics, including l2 (Euclidean distance), l1 (Manhattan distance), and dot product(Lin2019).
@@
SQ (Scalar Quantization): This method quantizes vector data to reduce storage and computation requirements(Johnson2017).
@@
NSG (Navigable Small World Graph): Another graph-based Approximate Nearest Neighbor (ANN) search method, it constructs a navigable small-world graph for fast retrieval(Fu2019).
LSH (Locality-Sensitive Hashing): This technique accelerates similarity search by mapping similar data into the same hash buckets(Datar2004).
@@ 2.3. LLMS Function Calling and Retrieval-Augmented Generation
LLM Function Calling enables large language models to interact with external functions, perform specific tasks, and generate more practical outputs(Gao2023). Function calling allows large language models to call external functions or services during the text generation process. By leveraging function calling, large language models can decide whether to invoke the corresponding functions based on instructions and return information in a structured format.
Retrieval-augmented generation (RAG) is a method that combines information retrieval and generation models to improve the performance of language models in handling complex queries and generation tasks. The core of the RAG method lies in its ability to combine pre-trained parametric memory with non-parametric external knowledge bases, accessing these external knowledge sources through a differentiable retrieval mechanism. This combination allows RAG models to rely not only on the internally learned knowledge during language generation but also to dynamically incorporate external information, thereby enhancing the relevance and accuracy of the generated content and eliminating model hallucinations(Lewis2020).
@@ 3.1. Dataset construction
Additionally, we needed a knowledge source database based on Retrieval-Augmented Generation (RAG) technology to mitigate hallucinations in large-scale language models while retaining crucial information. For this purpose, we employed the open-source large language model Mistral-7B-Instruct-v0.3(Mistral7BInstruct), which supports function calling. We designed precise prompts and detailed information extraction requirements for the model, embedding them into a function-calling framework. This enabled the large language model to automatically extract required geographic and historical information from Ctrip’s tourism information. Extracted information was formatted into a dictionary data structure, containing two primary key-value pairs: one for storing historical information and another for storing geographic information. Ultimately, we successfully extracted and organized the necessary geographic and historical information from the 563 viewpoints corresponding to our standardized database.
@@ 3.2. Database Vectorization and Indexing
The formula presented below is the formula for TF-IDF(Jones1972).
@@ 3.3. RAG Optimization for Large Language Models
We employed vector database retrieval to address the hallucination problem of large models. After identifying the viewpoints based on user needs, the geographical and historical information from the knowledge source database was returned to the large language model as an external knowledge base. Under this framework, the large language model no longer relies solely on its internal knowledge base to generate viewpoint introduction information. Instead, it understands and integrates the viewpoint introduction content through the external knowledge base. This method not only effectively optimize the hallucination problem of LLMs but also allows the model to fetch and generate comprehensive viewpoint introduction information, including geographical and historical details, even when certain knowledge is not included within the LLM itself. This research demonstrates the potential of RAG technology in enhancing the accuracy of information generated by large language models, providing new methods and tools for the standardization of tourism information and data analysis.
@@ 3.4. Evaluation Method
To evaluate the vector databases, we used the keyword-matching function of the SpaCy library along with a mature Chinese pre-trained model. After loading SpaCy’s(ExplosionAI) pre-trained Chinese model, we applied it to the result text data from different retrieval methods. By calculating the number of keywords that match the features in the prompts, we can determine the relevance and accuracy of the retrieval results.
Since the dataset used in this experiment is unstructured, we utilized the ”calculate composite score” scoring system’s b formula to calculate the scores for various dataset metrics and derive a composite score. Detailed scoring criteria can be referenced from Qi’s scoring standards(Qi2024).
@@
For the b formula, we rated the responses of the large language model from three aspects: fluency, accuracy rate, and relevance, and calculated the composite score. The weight coefficients for the different evaluation metrics, d1, d2, and d3, were set with a focus on solving the hallucination problem of the large model. We assigned relevance as the primary analysis metric, with weight parameters set to 0.3, 0.2, and 0.4, respectively. To ensure objectivity in scoring, we chose the fine-tuned Llama3-Chinese model(Wang2024) based on Wang’s work to evaluate the data results of each model. By analyzing the fluency of the generated sentences and the relevance of the generated text to the corresponding knowledge source content, we obtained the scores for fluency and relevance. We used the BERTscore(Zhang2019) evaluation tool developed by Zhang to score the generated text for the accuracy rate. By setting the prompt to control the scores within a range of 0 to 1, with 1 being the highest, we then applied the b formula to calculate the composite score for each model.
@@ 4.2. Optimization of Large Language Models with RAG Technology
In our RAG process, we utilized the LLaMA-Factory(Zheng2024), developed by hiyouga, as the tool for integrating knowledge and making inferences using large language models. We employed this tool with specific parameters to ensure the quality of the language model: a default temperature coefficient of 0.95, a maximum input length of 1024 tokens, and a maximum output length of 512 tokens. Below is the designed instruction for reference: ”Please use the provided viewpoint knowledge to introduce the viewpoint to the user. Always adhere strictly to the provided viewpoint information. Input: query of user, viewpoint information.”
@@
From the table, we can see that the performance of most experimental models improved after optimization with RAG technology. Keeping the weight parameters consistent and setting the scores of each metric to 1, we obtained a full score of 1.5873 using the formula. The ChatGLM3-6b(huggingface_chatglm3_2023) model achieved a total score of 1.2868 after fine-tuning, and a total score of 1.3228 after RAG optimization, with relevance increasing from 0.7885 to 0.7995, indicating improved model performance. The Baichuan2-7b(huggingface_baichuan2_2023) model had a total score of 1.0012 after fine-tuning, and a total score of 0.9215 after RAG optimization, with relevance decreasing from 0.5485 to 0.4687, suggesting that the model did not perform well after RAG optimization. This issue is likely due to the model’s poor performance, making it unsuitable for complex optimization techniques like RAG. The QWEN7b-chat(qwen2_2024) model achieved a total score of 1.1876 after fine-tuning and a total score of 1.3238 after RAG optimization, with relevance increasing from 0.6924 to 0.7966, showing a significant improvement and indicating that RAG technology effectively addresses the hallucination problem in this model. The Llama3-8b model(huggingface_meta_llama_2024) showed the most significant improvement, with relevance increasing from 0.6743 to 0.9721, making the model’s performance nearly perfect after optimization. We speculate that the significant impact is due to Llama3-8b’s lack of Chinese corpus, preventing effective integration and linking of content from the external knowledge base, resulting in overfitting.

### v2

@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: 10.48550/arXiv.2408.12003conference: September 20-22, 2024; Xiamen, China; isbn: 979-8-4007-1717-8submissionid: XA5000ccs: Computing methodologies Natural language generationccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processing
@@ 1.1. Current issues regarding Tibet’s cultural tourism and LLMs
To address these challenges, this study, under Project ID: XZ2024-01ZY0008, titled ”Research and Application of Intelligent Tourism Service System in Tibet Based on LLM,” aims to enhance tourists’ travel experiences and promote the development of Tibet’s tourism industry. The project focuses on the systematic organization and classification of viewpoints in the region. This includes detailed documentation of each viewpoint’s historical background, geographical location, cultural significance, and transportation information (Chen2020, ).
@@
Regarding the insufficiency in personalized recommendation capabilities, traditional models emphasize generalization but lack targeted fine-tuning, leading to suboptimal performance in personalized tourism contexts. Specifically, these models may exhibit ”hallucination” phenomena(Chen2023, ), where they fail to accurately understand user needs and recommend the most suitable attractions. This limitation is particularly evident in the context of Tibetan tourism, where the unique characteristics of many emerging attractions are not effectively captured, resulting in recommendations that lack accuracy and relevance.
Our research addresses these challenges by employing a vector database query method within the Retrieval-Augmented Generation (RAG) framework(meta_rag_2020, ). LLM extracts user requirements, which are then matched against a vector database to identify suitable attractions. These viewpoints are returned to the LLM for integration, resulting in accurate and contextually relevant recommendations tailored to the user’s preferences.
@@ 2.1. Vectorization Methods: TF-IDF and BERT
TF-IDF (Term Frequency-Inverse Document Frequency) is used to evaluate the importance of a word in a document collection or corpus. By multiplying TF (term frequency) and IDF (inverse document frequency), the TF-IDF value is obtained, which measures the significance of a word to a particular document(Havrlant2017, ).
BERT (Bidirectional Encoder Representations from Transformers) handles unlabelled text data by pre-training deep bidirectional representations(Devlin2019, ). This pre-training process involves jointly conditioning on both left and right context in all layers, allowing the pre-trained BERT model to be fine-tuned for various tasks by simply adding an output layer. The advantage of this approach lies in its flexibility to adapt to the text data of this project and its capability to create state-of-the-art models for a wide range of NLP tasks.
@@ 2.2. Retrieval Construction Methods
Flat (Brute Force Search): The system compares each vector in the database with the query vector, calculates their distance or similarity, and returns the closest vectors(Li2020, ).
HNSW (Hierarchical Navigable Small World): This method constructs a layered, small-world graph structure to accelerate query speed. HNSW supports multiple distance metrics, including l2 (Euclidean distance), l1 (Manhattan distance), and dot product(Lin2019, ).
@@
SQ (Scalar Quantization): This method quantizes vector data to reduce storage and computation requirements(Johnson2017, ).
@@
NSG (Navigable Small World Graph): Another graph-based Approximate Nearest Neighbor (ANN) search method, it constructs a navigable small-world graph for fast retrieval(Fu2019, ).
LSH (Locality-Sensitive Hashing): This technique accelerates similarity search by mapping similar data into the same hash buckets(Datar2004, ).
@@ 2.3. LLMS Function Calling and Retrieval-Augmented Generation
LLM Function Calling enables large language models to interact with external functions, perform specific tasks, and generate more practical outputs(Gao2023, ). Function calling allows large language models to call external functions or services during the text generation process. By leveraging function calling, large language models can decide whether to invoke the corresponding functions based on instructions and return information in a structured format.
Retrieval-augmented generation (RAG) is a method that combines information retrieval and generation models to improve the performance of language models in handling complex queries and generation tasks. The core of the RAG method lies in its ability to combine pre-trained parametric memory with non-parametric external knowledge bases, accessing these external knowledge sources through a differentiable retrieval mechanism. This combination allows RAG models to rely not only on the internally learned knowledge during language generation but also to dynamically incorporate external information, thereby enhancing the relevance and accuracy of the generated content and eliminating model hallucinations(Lewis2020, ).
@@ 3.1. Dataset construction
Additionally, we needed a knowledge source database based on Retrieval-Augmented Generation (RAG) technology to mitigate hallucinations in large-scale language models while retaining crucial information. For this purpose, we employed the open-source large language model Mistral-7B-Instruct-v0.3(Mistral7BInstruct, ), which supports function calling. We designed precise prompts and detailed information extraction requirements for the model, embedding them into a function-calling framework. This enabled the large language model to automatically extract required geographic and historical information from Ctrip’s tourism information. Extracted information was formatted into a dictionary data structure, containing two primary key-value pairs: one for storing historical information and another for storing geographic information. Ultimately, we successfully extracted and organized the necessary geographic and historical information from the 563 viewpoints corresponding to our standardized database.
@@ 3.2. Database Vectorization and Indexing
The formula presented below is the formula for TF-IDF(Jones1972, ).
@@ 3.3. RAG Optimization for Large Language Models
We employed vector database retrieval to address the hallucination problem of large models. After identifying the viewpoints based on user needs, the geographical and historical information from the knowledge source database was returned to the large language model as an external knowledge base. Under this framework, the large language model no longer relies solely on its internal knowledge base to generate viewpoint introduction information. Instead, it understands and integrates the viewpoint introduction content through the external knowledge base. This method not only effectively optimizes the hallucination problem of LLMs but also allows the model to fetch and generate comprehensive viewpoint introduction information, including geographical and historical details, even when certain knowledge is not included within the LLM itself. This research demonstrates the potential of RAG technology in enhancing the accuracy of information generated by large language models, providing new methods and tools for the standardization of tourism information and data analysis.
@@ 3.4. Evaluation Method
To evaluate the vector databases, we used the keyword-matching function of the SpaCy library along with a mature Chinese pre-trained model. After loading SpaCy’s(ExplosionAI, ) pre-trained Chinese model, we applied it to the result text data from different retrieval methods. By calculating the number of keywords that match the features in the prompts, we can determine the relevance and accuracy of the retrieval results.
Since the dataset used in this experiment is unstructured, we utilized the ”calculate composite score” scoring system’s b formula to calculate the scores for various dataset metrics and derive a composite score. Detailed scoring criteria can be referenced from Qi’s scoring standards(Qi2024, ).
@@
For the b formula, we rated the responses of the large language model from three aspects: fluency, accuracy rate, and relevance, and calculated the composite score. The weight coefficients for the different evaluation metrics, d1, d2, and d3, were set with a focus on solving the hallucination problem of the large model. We assigned relevance as the primary analysis metric, with weight parameters set to 0.3, 0.2, and 0.4, respectively. To ensure objectivity in scoring, we chose the fine-tuned Llama3-Chinese model(Wang2024, ) based on Wang’s work to evaluate the data results of each model. By analyzing the fluency of the generated sentences and the relevance of the generated text to the corresponding knowledge source content, we obtained the scores for fluency and relevance. We used the BERTscore(Zhang2019, ) evaluation tool developed by Zhang to score the generated text for the accuracy rate. By setting the prompt to control the scores within a range of 0 to 1, with 1 being the highest, we then applied the b formula to calculate the composite score for each model.
@@ 4.2. Optimization of Large Language Models with RAG Technology
In our RAG process, we utilized the LLaMA-Factory(Zheng2024, ), developed by hiyouga, as the tool for integrating knowledge and making inferences using large language models. We employed this tool with specific parameters to ensure the quality of the language model: a default temperature coefficient of 0.95, a maximum input length of 1024 tokens, and a maximum output length of 512 tokens. Below is the designed instruction for reference: ”Please use the provided viewpoint knowledge to introduce the viewpoint to the user. Always adhere strictly to the provided viewpoint information. Input: query of user, viewpoint information.”
@@
From the table, we can see that the performance of most experimental models improved after optimization with RAG technology. Keeping the weight parameters consistent and setting the scores of each metric to 1, we obtained a full score of 1.5873 using the formula. The ChatGLM3-6b(huggingface_chatglm3_2023, ) model achieved a total score of 1.2868 after fine-tuning, and a total score of 1.3228 after RAG optimization, with relevance increasing from 0.7885 to 0.7995, indicating improved model performance. The Baichuan2-7b(huggingface_baichuan2_2023, ) model had a total score of 1.0012 after fine-tuning, and a total score of 0.9215 after RAG optimization, with relevance decreasing from 0.5485 to 0.4687, suggesting that the model did not perform well after RAG optimization. This issue is likely due to the model’s poor performance, making it unsuitable for complex optimization techniques like RAG. The QWEN7b-chat(qwen2_2024, ) model achieved a total score of 1.1876 after fine-tuning and a total score of 1.3238 after RAG optimization, with relevance increasing from 0.6924 to 0.7966, showing a significant improvement and indicating that RAG technology effectively addresses the hallucination problem in this model. The Llama3-8b model(huggingface_meta_llama_2024, ) showed the most significant improvement, with relevance increasing from 0.6743 to 0.9721, making the model’s performance nearly perfect after optimization. We speculate that the significant impact is due to Llama3-8b’s lack of Chinese corpus, preventing effective integration and linking of content from the external knowledge base, resulting in overfitting.